40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Mon Aug 10 02:11:47 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 IMG_20150803_223403975.jpgIMG_20150803_221210816_HDR.jpgIMG_20150803_223420267.jpg
    Reply  Mon Aug 10 11:39:40 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 
       Reply  Mon Aug 10 11:57:17 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 screen_shot.png
          Reply  Mon Aug 10 12:09:49 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 IMG_20150809_215628585_HDR.jpg
             Reply  Fri Aug 28 01:08:14 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ Calibration RFAM_detector_calib.pdfRFAM_detector_calib_spectra.pdf
                Reply  Fri Aug 28 02:14:53 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ 37MHz OCXO AM measurement OCXO_AM_noise.pdf
                   Reply  Tue Sep 8 10:55:31 2015, Koji, Electronics, AM Stabilized EOM Driver, RF AM Measurement Unit E1500151 ~ 37MHz OCXO AM measurement 
Message ID: 232     Entry time: Mon Aug 10 11:39:40 2015     In reply to: 231     Reply to this: 233
Author: Koji 
Type: Electronics 
Category: AM Stabilized EOM Driver 
Subject: RF AM Measurement Unit E1500151 

Entry for Aug 6th, 2015

I faced with difficulties to operate the RF AM detectors.

I tried to operate the RF AM detector. In short I could not as I could not remove the saturation of the MON outputs, no matter how jiggle the power select rotary switches. The input power was 10~15dBm.

D0900761 Rev.A
https://dcc.ligo.org/LIGO-D0900761-v1

I've measured the bias voltage at TP1. Is the bias such high? And does it show this inversion of the slope at high dBm settings?

Setting Vbias
 [dBm]   [V]

   0    21.4
   2    21.0
   4    20.5
   6    19.8
   8    18.8
  10    17.7
  12    16.3
  14    14.2
  16    11.9
  18    10.6
  20    11.8
  22    15.0


The SURF report (https://dcc.ligo.org/LIGO-T1000574) shows monotonic dependence of Vbias from 0.6V to 10V (That is supposed to be the half of the voltage at TP1). 
I wonder I need to reprogram FPGA?

But if this is the issue, the second detector should still work as it has the internal loop to adjust the bias by itself.
TP3 (Page 1 of D0900761 Rev.A) was railed. But still MON2 was saturated.
I didn't see TP2 was also railed. It was ~1V (not sure any more about the polarity). But TP2 should also railed.

Needs further investigation

ELOG V3.1.3-